EXCEPTIONAL THETA-CORRESPONDENCES I

Kay Magaard Gordan Savin

Abstract. Let G be a split simply laced group defined over a p-adic field F. In this paper we study the restriction of the minimal representation of G to various dual pairs in G. For example, the restriction of the minimal representation of E_{7} to the dual pair G_{2} × Sp(6) gives the non-endoscopic Langlands lift of irreducible representations of G_{2} to Sp(6).

Typeset by AMS-TEX